OpenStudy (anonymous):

find the values of the constants a and b such that lim x goes to 0 (square root)a+bx-(square root)3/x=(square root)3

OpenStudy (ash2326):

\[\lim_{x\to 0} \frac{\sqrt{a+bx}-\sqrt 3}{x}=\sqrt 3\] Is this your question?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

For the limit to exist, when x goes to 0 Both numerator and denominator should become 0, denominator becomes 0 when x=0 could you equate the value of numerator to zero, plugging in x=0??

OpenStudy (anonymous):

no i do not know how to do that

OpenStudy (anonymous):

is it just the square root of a+b - the square root of 3?

OpenStudy (ash2326):

We have \[\sqrt{b+ax}-\sqrt {3}=0. For\ x=0\] put x=0 here \[\sqrt b-\sqrt 3=0\] what do you think is the value of b?

OpenStudy (anonymous):



OpenStudy (ash2326):

That's root b, value of b?

OpenStudy (anonymous):

9

OpenStudy (ash2326):

if we had \[\sqrt b=3\longrightarrow b=9\] \[\sqrt b=\sqrt 3\longrightarrow b=???\]

OpenStudy (anonymous):

\[\sqrt{9} ???\]

OpenStudy (ash2326):

square both sides, what'd you get for b?